Hot water cools from 60degree celcius to 50 degree celcius in the first 10 min and to 42 degree celcius in the next 10 min .The temp of the surrounding.

Grade:12

3 Answers

Ramesh V
14 years ago

average temp. in first case = (60+50) /2 = 55 C

average temp. in second case = (50+42) /2 = 51 C

as frm law of cooling

dT/dt = -k(T-To )  where To is temp. of surrounding

case 1 : dT = 10 C and dt = 10 mins

so ,  1=-k(55-To )         ....(1)

case 2 : dT = 8 C and dt = 10 mins

so ,  0.8=-k(51-To )       .........(2)

From 1 and 2 , we have

Temp of surr. as 35 C

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch

 

 

Visalakshi Senthil Kumar
8 years ago
Excuse sir I need to correct u in the second step ie., (50+42)/2=46 
So the room temp would be 10C    
dT/dt=0.8=-k (46-t)
dT2/dt2 =k (55-t)
Dividing them both we get
1/0.8=(55-t)/(46-t)
Therfore t=10
Answer by Aswath S, AECS PU College, Bangalore Fiitjee
Future IIT ian     ha ha ha
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
Average temp. in first case = (60+50) /2 = 55 C
Average temp. in second case = (50 + 42) / 2 = 46 C
from law of cooling
dT/dt = – k(T – To )  where To is temp. of surrounding
 
Case 1 : dT = 10 C and dt = 10 mins
so ,  1 = – k(55 – To)
Case 2 : dT/dt = 0.8 = – k (46 – To)
0.8 = k(55 – To)
Dividing them both we get
1/0.8 = (55 – To) / (46 – To)
On solving, we get, To = 10oC
 
Hope it helps.
Thanks and regards,
Kushagra